### v3

# TrustSQL: Benchmarking Text-to-SQL Reliability with Penalty-Based Scoring

Abstract
Text-to-SQL enables users to interact with databases using natural language, simplifying the retrieval and synthesis of information. Despite the remarkable success of large language models (LLMs) in translating natural language questions into SQL queries, widespread deployment remains limited due to two primary challenges. First, the effective use of text-to-SQL models depends on users’ understanding of the model’s capabilities—the scope of questions the model can correctly answer. Second, the absence of abstention mechanisms can lead to incorrect SQL generation going unnoticed, thereby undermining trust in the model’s output. To enable wider deployment, it is crucial to address these challenges in model design and enhance model evaluation to build trust in the model’s output. To this end, we introduce TrustSQL, a novel comprehensive benchmark designed to evaluate  text-to-SQL reliability—defined as a model’s ability to correctly handle any type of input question by generating correct SQL queries for feasible questions and abstaining from generating infeasible ones (e.g., due to schema incompatibility or functionalities beyond SQL). We evaluate existing methods using a novel penalty-based scoring metric with two modeling approaches: (1) pipeline-based methods combining SQL generators with infeasible question detectors and SQL error detectors for abstention; and (2) unified methods using a single model for the entire task. Our experimental results reveal that achieving high scores under severe penalties requires significant effort and provide a new perspective on developing text-to-SQL models for safer deployment.
1 Introduction
Text-to-SQL is a task that involves generating SQL queries from natural language (NL) questions, a challenge that has persisted for several decades shwartz2022tabular . This task is crucial both within and beyond the realm of machine learning and natural language processing (NLP), due to the ubiquity of databases in various fields such as medicine, finance, retail, and telecommunications. Accessing structured data stored in these databases through natural language can unlock the true potential of big data, enabling domain experts to uncover new quantitative insights van2024tabular .
To effectively utilize these models, users need to be familiar with their capabilities. These include the types of questions each individual model can accurately answer, the information it has access to, and what operations can or cannot be performed via SQL. However, many potential users who could greatly benefit from the model are unfamiliar with these capabilities and may find it challenging to use, as their questions may frequently fall outside the model’s scope (beyond the correct SQL in Figure 1(b)). The challenge intensifies in real-world deployment scenarios where user questions include not only those feasible for the model but also infeasible ones, such as requests requiring knowledge beyond the database content or SQL functionalities (as shown in Figure 1(a)). Moreover, the absence of abstention mechanisms in current text-to-SQL modeling can lead to incorrect model outputs for both feasible and infeasible questions going unnoticed, potentially outweighing the benefits of model use. If a model, by design, could selectively answer correct SQL queries while abstaining from answering the rest, it would significantly boost user confidence, thereby enhancing trust in the model and encouraging its use whitehead2022reliable ; ren2022out .
Figure 1: (a) Deployment Scenarios include both feasible and infeasible user questions regarding text-to-SQL tasks (e.g., requests outside the database schema or SQL functionalities). (b) Standard Text-to-SQL aims to maximize correct SQL query generation. This focus can result in incorrect model outputs going unnoticed, increasing risk in deployment. The green and red regions represent correct and incorrect SQL generations, respectively. To effectively utilize a text-to-SQL model, users must be familiar with its capabilities, which vary by model—the types of questions each model can correctly answer—setting a high barrier for potential users. (c) Reliable Text-to-SQL removes this barrier and aims to enhance trust in the model by minimizing incorrect outputs (III) and (IV) through abstention (red arrow), while maximizing correct outputs (I) within the feasible region (green arrow). By shrinking the red regions (III) and (IV), and subsequently the yellow region (II), users will begin to trust the model’s outputs. Each region is further explained in the footnote222.
22footnotetext: The five key regions in Figure 1(c) are: (I) correct SQL generation for feasible questions (desirable for text-to-SQL reliability), (II) abstention for feasible questions (neither strictly desirable nor undesirable, thus neutral), (III) incorrect SQL generation for feasible questions (undesirable), (IV) incorrect SQL generation for infeasible questions (undesirable), and (V) correct abstention for infeasible questions (desirable). All five distinct cases are all incorporated into our penalty-based scoring metric introduced in Section 3.2, and their actual examples are provided in Supplementary E.
To address the aforementioned challenges, we present TrustSQL, a new benchmark designed to assess text-to-SQL reliability. We define text-to-SQL reliability as a model’s ability to handle any input questions correctly through accurate SQL generation and abstention, thereby reducing the red and yellow regions by covering the entire region in green, as shown in Figure 1(c)). Unlike standard text-to-SQL tasks defined within the feasible region (Figure 1(b)), real user questions span both feasible and infeasible task regions, including requests beyond the database content and SQL functionalities (Figure 1(a)). We constructed this new comprehensive setting using three domain-specific datasets—ATIS atis , Advising advising , and EHRSQL lee2022ehrsql —along with one cross-domain dataset—Spider yu2018spider . In constructing TrustSQL, we reflected common error cases undermining text-to-SQL reliability through careful data splitting and incorporation of various infeasible questions, similar to red teaming. We then evaluated existing methods using a novel penalty-based scoring metric designed specifically to measure text-to-SQL reliability (details in Section 3.2). Our experimental results reveal that while most advanced methods seem to perform well in penalty-less settings, achieving high scores under severe penalty settings remains highly challenging. From the high penalty perspective, knowing when to abstain precisely carries more weight than rashly attempting to generate SQL queries, underscoring the need for substantial efforts towards gaining trust in model outputs.
2 Dataset Construction
In this section, we begin by identifying common error cases that undermine text-to-SQL reliability during deployment, causing answers to reside in the red regions in Figure 1(c). Then, we introduce our strategies to incorporate each case into the dataset construction process. Below are the common error cases in text-to-SQL deployment:
  • Complexity of SQL generation: User questions, especially those posed by domain experts, are often too complex for models to answer accurately. This complexity arises from the inherent difficulty of the queries yu2018spider ; li2023can ; hazoom2021sede ; lee2022ehrsql , as in Table 6 in Supplementary B, or from the model encountering unfamiliar phrases and entities specific to the database suhr-etal-2020-exploring ; lee-etal-2021-kaggledbqa . Models that generate SQL without the ability to abstain from such cases risk compromising trust in their output. We include these dimensions of query difficulty [P1] and familiarity [P2] in feasible data construction.
  • Ambiguity in user queries: User questions are often ambiguous, particularly when they refer to columns or values (i.e., database records) without precise specification. Resolving the user’s true intent involves multi-turn interactions for clarification, which is beyond the scope of this work.
  • Schema incompatibility: Users who are unfamiliar with the exact database schema may inadvertently ask questions that cannot be processed because the information they seek does not exist in the existing columns. Previous works construct such cases by deleting and perturbing existing columns triagesql ; known_dte . Instead, we include three distinct types within this case [P3] to provide more granular views on model behavior.
  • Scope of SQL functionalities: Users may not always be familiar with the scope of SQL functionalities, leading to requests that require complex numerical operations, data analysis, or scheduling appointments, which are beyond standard SQL execution. We include such questions [P4] in infeasible data construction.
  • Need for external knowledge: Users may ask questions that involve factual world knowledge that goes far beyond the database content, either fully (e.g., ‘What airport is in Wilder, KY?’) or partially (e.g., ‘Find me patients born after WWII’). We consider the former cases [P5] when constructing infeasible questions, while excluding the latter.
Table 1: Sample infeasible questions for each type for the orchestra database in Spider, containing 4 tables with 23 columns. Samples for domain-specific databases are not shown here for brevity. ‘Note’ indicates hypothetical columns in existing tables (i.e., columns not present in the database we use for text-to-SQL) or a task name used for question annotation.
Type Question Note column-surface What is the average attendance capacity for all the shows? show.attendance_capacity column-related Which orchestra has the most members? orchestra.number_of_member column-unrelated What aircraft does Charles Dutoit own? conductor.aircraft non-sql Identify the top three significant features that influence attendance. feature importance ext-know What advantages does diversity bring to an orchestra? -
2.1 Feasible Data Construction
Dataset selection
To establish the benchmark data covering comprehensive settings, we selected three complex domain-specific text-to-SQL datasets, namely, ATIS (flight information) atis , Advising (education) advising , and EHRSQL (healthcare) (lee2022ehrsql, ). Additionally, we included one cross-domain dataset, Spider. The distinction between domain-specific and cross-domain text-to-SQL settings, along with the complexity of each dataset, is detailed in Supplementary B.
Data preprocessing
Our preliminary data analysis revealed that ATIS and Advising require significant preprocessing due to issues such as numerous query duplications, inconsistencies, and poor-quality natural language (NL) questions (each issue discussed in SupplementaryG). We implemented several strategies to correct these issues, including creating new question paraphrases for all input NL questions using GPT-4, followed by a manual review by the authors. For further details, please refer to Supplementary G). The databases of the four datasets and their corresponding question-SQL pairs, after preprocessing, are included as questions in the feasible task region (light blue region in Figure 1(a)). To reflect the query difficulty [P1], we adopted the SQL hardness criteria following pourreza2023din-sql across all datasets. Detailed statistics for each hardness level are provided in Table 6 in Supplementary B. Query familiarity [P2] is later introduced in the data splitting process (§2.3).
Table 2: Dataset statistics of TrustSQL. TrustSQL includes an equal number of unanswerable and answerable questions in the test sets. For $Q$, the three numbers, separated by ’/’, denote the size of training, validation, and test sets.
ATIS Advising EHRQSL Spider
All Questions ($Q\cup Q^{c}$) 2555 2769 7473 8561
   Feasible Question ($Q$) 1114 / 489 / 476 1170 / 533 / 533 4674 / 931 / 934 7000 / 507 / 527
   Infeasible Question ($Q^{c}$) 476 533 934 527
    non-sql 99 110 191 100
    surface 93 102 185 100
    unrelated 94 110 187 100
    related 98 107 189 100
    ext-know 182 127 92 104
2.2 Infeasible Data Construction
Text-to-SQL models deployed in real-world scenarios often encounter user requests that fall into the infeasible task region (gray region in Figure 1(a)) due to going beyond the database content or SQL functionalities, and thus do not have corresponding SQL queries. To evaluate models in a more realistic environment, we manually annotated NL questions by incorporating three common infeasible error cases as follows:
Schema incompatibility
[P3]: Instead of randomly dropping existing columns in the database triagesql ; known_dte , we introduce the concept of hypothetical columns as target columns for infeasible questions. Those questions are asked as if their corresponding SQL queries exist, but in fact do not, since no such columns exist in the database. We categorized these into three types: column-surface, column-related, and column-unrelated, representing attacks based on surface-form and semantic column similarities and dissimilarities, respectively.
  • column-surface: Questions referring to hypothetical columns that resemble existing columns in different tables, or have similar surface forms (e.g., a hypothetical ‘age’ column in the ‘Professionals’ table that resembles the ‘age’ column in the ‘Dogs’ table in the ‘dog_kennels’ database).
  • column-related: Questions referring to hypothetical columns that could naturally integrate into the existing database schema (e.g., a hypothetical column named ‘duration’ in the ‘song’ table). This question type can be seen as a manual version of the ADD perturbations in ADVETA adveta .
  • column-unrelated: Questions referring to hypothetical columns that are randomly sampled from other databases, which are mostly unnatural (e.g., a hypothetical ‘product_category’ column in the ‘battles’ database).
Beyond SQL functionalities
[P4]: To capture diverse user requests that exceed SQL execution (non-sql), we compiled a list of NLP and data analytics tasks and used GPT-4 to identify what could be done as an AI assistant if any tools can be used given the database content. Based on a list of 27 tasks (e.g., anomaly detection, visualization, and web search), we manually annotated NL questions relevant to or referring to existing columns in each database (thus, context-driven annotation), simulating users expecting the model to perform tasks not solely addressed through SQL.
Need for external knowledge
[P5]: To construct NL questions involving external knowledge (ext-know), we prompted GPT-4 to generate questions that require domain expertise, qualitative insights, or causal understanding. We continued this process until we achieved a balanced set of samples comparable to other types of infeasible questions.
After manual annotation of infeasible NL questions, the authors used GPT-4 to iteratively review them to ensure that they were truly infeasible. The sample annotated questions are provided in Table 1.
2.3 Data Splitting
The core principle for data splitting is to create varying query familiarity [P2] for feasible data and introduce the unexpectedness of infeasible data for model evaluation. Table 2 summarizes the overall data statistics in TrustSQL.
Feasible data
We define ‘seen SQL’ and ‘unseen SQL’ to introduce query familiarity [P2] (only applicable within feasible data). Samples are considered ‘seen SQL’ when their corresponding SQL structures [^3] A SQL structure refers to a specific combination of SQL components with specific tables and columns, excluding actual data values. are present in the training set but have different paraphrased NL questions in the validation and test sets. ‘Unseen SQL’ samples have SQL structures not present in the training set. We use these different query familiarities [P2] included in the validation and test sets to analyze model performances. The data splitting process for domain-specific datasets is as follows: 1) Split feasible data into training, validation, and test sets in a 6:2:2 ratio based on NL questions (same SQL structures across splits, but different NL questions); 2) Remove 20% of non-overlapping SQL structures from the validation and test sets; 3) Remove the total 40% of these SQL structures from the training set, resulting in 60% ‘seen SQL’ in all splits, while validation and test sets contain 80% SQL structures (60% ‘seen SQL’ and 20% ‘unseen SQL’). For Spider, we divide the development split in half for new validation and test sets (10 databases each), all ‘unseen SQL,’ with no overlaps with the training set.
Infeasible data
We add the annotated infeasible NL questions of five types (column-surface, column-related, column-unrelated, non-sql, and ext-know) all to the test set in balanced proportions to reflect their unexpected encounter in deployment.
3 Benchmarking Text-to-SQL Reliability
3.1 Task
Reliable text-to-SQL aims to develop models that only generate correct SQL queries while abstaining in any other cases, including both feasible ($\mathcal{Q}$) and infeasible questions ($\mathcal{Q}^{c}$), to enhance text-to-SQL reliability. Optimized under this principle, users can effectively use these models without having extensive knowledge about the models’ capabilities. They can also be less concerned about the model producing incorrect outputs, as it would abstain from answering questions beyond its capabilities. This increased reliability will lead to wider adoption of text-to-SQL models, empowering domain experts to utilize natural language interfaces to access and analyze the vast amounts of data stored in databases.
3.2 Evaluation Metric
In evaluating reliable text-to-SQL performance, we categorize the model’s outputs into three distinct categories: desirable (green), undesirable (red), and neutral (yellow), as illustrated in Figure 1(c). These categories reflect different types of consequences brought about by the model’s outputs. Desirable outputs, providing utility, include correct SQL generation (I) and abstentions (V). However, the severity of the consequences of undesirable outputs can vary based on the context of the deployment. In less critical settings (e.g., experimental scenarios), we might be lenient towards these undesirable outputs and focus solely on the model’s utility. In contrast, in high-stakes environments (e.g., hospitals), undesirable outputs can have severe repercussions, potentially outweighing any utility provided by the model. Therefore, we define the relative magnitude of the negative impact compared to the positive impact as risk. This risk includes incorrect SQL generations (III) and abstentions (IV), weighted by a penalty factor, $c$, namely the ratio of the negative impact to the positive impact. Neutral (yellow) cases (II) are neither strictly desirable nor strictly undesirable; therefore, we consider them neither utility nor risk. Examples of instances for each case are provided in Appendix E.
In summary, text-to-SQL reliability can be quantified as the difference between utility and risk. To formalize this notion mathematically, we define this penalty-based scoring for text-to-SQL—reliability score (RS)—as follows:
$\phi_{c}(x)=\begin{cases}1&\text{if $x\in\mathcal{Q}$; $g(x)=1$; $Acc(x)=1$,}\\ 0&\text{if $x\in\mathcal{Q}$; $g(x)=0$,}\\ -c&\text{if $x\in\mathcal{Q}$; $g(x)=1$; $Acc(x)=0$,}\\ -c&\text{if $x\in\mathcal{Q}^{c}$; $g(x)=1$,}\\ 1&\text{if $x\in\mathcal{Q}^{c}$; $g(x)=0$,}\\ \end{cases}$ (1)
where $Acc(x)$ is the execution accuracy in text-to-SQL and $g(x)=1$ and $g(x)=0$ indicate the model’s decision to answer or abstain, respectively.
For each sample, a score of 1 is assigned when the SQL is generated correctly for $x\in\mathcal{Q}$ (I) or when the model abstains for $x\in\mathcal{Q}^{c}$ (V), indicating utility. A score of 00 is given when the model decides not to generate an SQL for $x\in\mathcal{Q}$ (II), indicating no risk and no utility. A penalty of $-c$ is assigned when the generated SQL is incorrect (III) or when the model attempts to generate an SQL for $x\in\mathcal{Q}^{c}$ (IV), indicating risk. Actual examples of each case are provided in Supplementary E. After calculating each sample-level score, the aggregate RS can be computed by taking the mean, $\frac{1}{N}\sum_{x}\phi_{c}(x)$, resulting in $\phi_{c}$.
In our experiments, we consider three options for the penalty value $c$: 0, 10, and $N$, where $N$ represents the total number of samples to be evaluated.
  • $c=0$: The most lenient scenario. This does not penalize any mistakes made by the model. If we only handle feasible questions, this score is equivalent to the standard text-to-SQL accuracy.
  • $c=10$: A moderately strict scenario. Every 10 correct answers weigh the same as one incorrect answer.
  • $c=N$: The most strict scenario. A single mistake, even if all other answers were correct, leads to a negative score.
An RS($c$) of 100% indicates perfect text-to-SQL reliability, meaning the model can handle all inputs correctly (i.e., covering all the regions with green in Figure 1(c)). RS($c$) > 0 means the utility is greater than the risk, making it deployable in the context of text-to-SQL reliability.
4 Experiments
4.1 Modeling
In this section, we describe various methods for developing models to enhance text-to-SQL reliability. We categorize the methods into two broad modeling approaches: pipeline-based and unified. For each approach, we develop both open-source and commercial API-based models, as these choices are often utilized one or the other depending on various factors in deployment, such as budget constraints, technical expertise, and privacy concerns.
Table 3: Model comparison in TrustSQL involving both feasible and infeasible questions evaluated in $\phi_{{}_{0}}$, $\phi_{{}_{10}}$, and $\phi_{{}_{N}}$. AT, AD, EH, and SP refer to ATIS, Advising, EHRSQL, and Spider, respectively. Abstain-All blindly abstains from predicting all input questions, reflecting the ratio of infeasible questions. All in % for readability.
[…]
4.1.1 Pipeline-based Approach
In conventional NLP task formulations, enhancing text-to-SQL reliability can be addressed by combining three different tasks: infeasible question detection triagesql ; known_dte , SQL generation pourreza2023din-sql ; wang2023mac , and SQL error detection chen2023teaching .
Open-source
For SQL generation, we use T5 (T5-3B), feeding questions followed by a serialized schema listing all tables and columns suhr-etal-2020-exploring . To incorporate abstention mechanisms, we use SQLCoder (sqlcoder-7b-2) for infeasible question detection (before SQL generation) and error detection (after SQL generation). For infeasible question detection, we fine-tune SQLCoder on TriageSQL triagesql , denoted as CLSM. For error detection, we train SQLCoder based on the generated SQLs and their sample-level accuracy from trained T5 on the TrustSQL validation set plus samples of infeasible questions in TriageSQL, denoted as ErrorM. Details about training these models are provided in Supplementary C. When denoting the combined methods, we use arrows to indicate data flow between models, such as CLSMT5 indicating that question detection comes first, followed by SQL generation.
API-based
For SQL generation, we employ three methods using GPT-3.5 (gpt-3.5-turbo-0125): SQLPrompt, SQL generation with few-shot prompting chang2023prompt , DIN-SQL pourreza2023din-sql , a state-of-the-art task decomposition framework in Spider, and MAC-SQL wang2023mac , a state-of-the-art multi-agent framework in BIRD li2023can . We adjust text prompts for DIN-SQL and MAC-SQL for domain-specific datasets. For abstention, we use GPT-4o for both infeasible question detection and error detection. We provide the modeling details in Supplementary D.
4.1.2 Unified Approach
To achieve text-to-SQL reliability using a single model, we explore various text-to-SQL prompts that incorporate abstention mechanisms. Additionally, we utilize multiple uncertainty estimation methods, leveraging SQL generators’ hidden representations and probabilities, to enable abstention.
Open-source
For SQL generation, we use T5 (T5-3B) due to its wide use in selective generation ren2022out and calibration stengel2023calibrated in NLP. For abstention, we leverage the hidden representations of the input and the output probabilities by adopting uncertainty estimation to detect samples that are out-of-distribution (OOD) for abstention. Below are the methods we use:
  • Entropy-based: Each sample is assigned the maximum entropy value among the generated SQL tokens, denoted as MaxEnt (lee2022ehrsql, ).
  • Probability-based: Each sample is assigned the minimum of the highest token probabilities from the token sequence, denoted as MaxProb (stengel2023calibrated, ).
  • Distance-based: Mahalanobis distance calculation based on the distribution of the training data representations, denoted as FeatMD (lee2018simple, ; ren2022out, ), and relative Mahalanobis distance calculation with a background distribution using the TriageSQL dataset, denoted as FeatRMD (ren2021simple, ; ren2022out, ).
These methods require calibrating thresholds to draw a decision boundary between in-distribution and out-of-distribution. Our proposed heuristic method is provided in Supplementary C.2.
Table 4: Model comparison in TrustSQL using only feasible data. In this table, $\phi_{{}_{0}}$ is equivalent to execution accuracy in standard text-to-SQL. The scores in shaded cells represent SQL performance without abstention.
[…]
API-based
We use two prompting methods with GPT-3.5, as explained below:
  • SQLPrompt[DEMO]: Few-shot demonstrations of similar feasible questions with their corresponding SQL queries and similar infeasible questions, each labeled with ‘not_answerable’ sun2021conditionalqa . The order of demonstrations is randomly chosen.
  • SQLPrompt[Voting]: Using the same prompts as SQLPrompt, we sample the answers five times (wang2022self, ) and consider the SQL generation only if all samples unanimously agree; otherwise, we abstain.
5 Results
5.1 Model Comparison
Full Data
Table 3 shows the performance of each method in the TrustSQL benchmark. For open-source models, the probability-based (i.e., T5[MaxEnt]) and entropy-based (i.e., T5[MaxProb]) methods in the unified approach consistently outperform other methods across different penalty settings. For API-based models, CLSPSQLPrompt performs best in RS(0), while CLSPSQLPromptErrorP excels in stricter settings (i.e., RS(10) and RS(N)). This indicates that both question detection and error detection are crucial for enhancing text-to-SQL reliability. SQLPrompt[Voting] shows notable performance across all penalties, particularly given its simplicity.
Feasible Data Only
Evaluating models on feasible data measures unnecessary abstention and the impact of error detectors, indicated by the performance drop between the SQL performance without abstention (shaded cells) and the models with abstention in RS(0). Similar to the full data setting, open-source models, T5[MaxEnt] and T5[MaxProb], tend to perform best with increasing penalties. API-based models using the unified approach, particularly SQLPrompt[Voting], outperform SQLPrompt[Demo] as the penalty increases. High performance in RS(0) indicates frequent abstention, resulting in lower performance with higher penalties. A more detailed analysis of each component in pipeline-based methods is provided in Supplementary F.
5.2 Discussion
Q1: Why is All-abstain considered deployable in text-to-SQL reliability?
In the context of standard text-to-SQL, All-abstain in Table 3 cannot be considered deployable since it contradicts the very purpose of the task. However, in the context of text-to-SQL reliability, All-abstain achieves an RS of 50 on all $c$ by correctly handling 50% of all input questions (i.e., the proportion of infeasible questions in TrustSQL) with certainty. This approach may be more reliable (absolutely no risk in all settings) than a model optimized for standard text-to-SQL, which always generates SQL queries that may be incorrect (risk potentially higher than utility).
Q2: What is an easy way to enhance text-to-SQL reliability?
An easy way to enhance the RS to at least a deployable level (i.e., >0) in strict settings is to generate only those SQL queries for which the model has high confidence, while abstaining from generating any when uncertain. The experimental results demonstrate that this approach is particularly effective in settings where incorrect generation carries a high penalty. For instance, as shown in Table 3, the T5[FeatMD] model on Spider generates only one correct SQL query and abstains from the rest due to our strict heuristic thresholding (see Supplementary C.2). Efforts to reduce risk can gain more RS scores than efforts to increase utility. However, achieving higher RS (i.e., >50%) with such cautious SQL generation requires another line of effort, which has not been explored extensively.
Q3: Why is TrustSQL necessary as a text-to-SQL benchmark?
Existing benchmarks lack penalty-based scoring and considerations for infeasible questions. TrustSQL addresses these gaps by modeling SQL generation with abstention and handling error cases from infeasible question detection, which are then propagated to SQL generation and error detection. This comprehensive approach encourages new modeling aspects to enhance reliability. For instance, the entries column-surface and non-sql in Table 1 are examples that fail to be handled using GPT-4o-based infeasible question detection and error detection methods (i.e., CLSPSQLPromptErrorP). Conventional task settings cannot explore such cases, showing how these models respond to infeasible questions. For instance, the query ‘SELECT avg(attendance) FROM show’ is generated for the question ‘What is the average attendance capacity for all the shows?’ (where only the attendance column exists, not the capacity), returning a result of 1326.4 from the database. TrustSQL’s comprehensive approach provides a new perspective on error cases in text-to-SQL generation, encouraging the development of more reliable models.
6 Related Works
Model Reliability in Text-to-SQL
Text-to-SQL has attracted significant attention for its practical applications. Key single-domain datasets include ATIS (atis, ), GeoQuery (geography, ), Advising advising , SEDE hazoom2021sede , and EHRSQL lee2022ehrsql . For cross-domain tasks, Spider yu2018spider and BIRD li2023can serve as essential benchmarks for SQL generation. However, these benchmarks assume that all input questions should generate SQL, lacking incentives for developing models to abstain. This absence of safety measures undermines reliable deployment. Some efforts focus on detecting infeasible questions given database schemas triagesql ; known_dte , but they do not integrate SQL generation or consider the error-handling capabilities of text-to-SQL models. Our work aims to integrate comprehensive reliability measures, aligning the benchmark more closely with model trust.
Uncertainty Estimation in NLP
Uncertainty estimation is vital for AI reliability hendrycks2021unsolved . The methods include probability-based methods (calculating entropies or model-assigned probabilities)marek2021oodgan ; xiao2021hallucination ; kim2022uncertainty ; stengel2023calibrated and distance-based methods (using metrics such as the Mahalanobis distance to compare new inputs with training data)lee2018simple ; zheng2020out ; ren2021simple ; ren2022out . Probability-based methods gauge model confidence, enhancing output reliability. Distance-based methods identify outliers likely to produce low-quality outputs, promoting safer AI deployment. Our work employs both approaches to develop abstention mechanisms for text-to-SQL models.
7 Conclusion and Future Direction
In this work, we present TrustSQL, a new benchmark designed to assess text-to-SQL reliability, defined as models’ ability to correctly handle any type of input question by generating correct SQL queries for feasible questions and abstaining from infeasible ones. To demonstrate this, we constructed the benchmark based on common error cases that users unfamiliar with the model’s capabilities may encounter. We developed and assessed two approaches: pipeline-based and unified methods, using existing modeling strategies. Our experimental results, using a penalty-based scoring metric, show that addressing this task requires various modeling approaches and opens new avenues for model development.
Though we have carefully designed the benchmark, there are several limitations. First, despite our efforts to cover a wide range of infeasible questions, our benchmark does not encompass all those found in real-world scenarios. Secondly, our work focuses on single-turn text-to-SQL tasks and does not address the complexities of disambiguating uncertainty in entity and schema linking or user intents. Finally, results of commercial API-based models in our experiment may not be reproducible if any updates are made.
Future research directions include developing agent-based methods that can plan ahead to achieve high performance in RS(N), bypassing the significant need for strategies such as prompt engineering and domain adaptation, accelerating wider adoption of text-to-SQL across diverse fields.